# MD 20742, USA

### v2

@@ 2. An analogue of LIL for Smooth expanding maps
(b) If $(M4)_{*}$ holds, then with probability $1$, $A^{n}_{*}$ does not occur for large $n.$
Suppose ${\mathbf{S}}_{r}=\infty.$ For any $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r}$ and any intervals $I_{1},I_{2}\dots I_{r}\subset[0,1]:$
If
(c) $(M1)_{k}$, $(M2)_{k}$ for $k=r+1,\dots,2r$ as well as $(M3)_{r}$ are satisfied and $\hat{s}(n)/n<\min_{0\leq i,\,j\leq r,\,\,i\neq j}d(I_{i},I_{j}),$ where $I_{0}=0;$ or
(d) $(M1)_{r},$ $(M3)_{r}$ are satisfied and $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$ holds for $\hat{s}(n)/n<\min_{0\leq i,\,j\leq r,\,\,i\neq j}d(I_{i},I_{j}),$ where $I_{0}=0,$ then with probability $1$, there are infinitely many $n$ such that for some $k_{1}(n),k_{2}(n)\dots k_{r}(n)$ with $k_{j}(n)/n\in I_{j},$ the events $\Omega^{k_{j}}_{\rho_{n},i_{j}}$ occur.
Proof.
@@
Part (a) is Theorem 2.6 in [12]. We consider the case ${\mathbf{S}}_{r}<\infty$ and $(M4)_{*}$ holds. Suppose $\{\delta_{n}\}$ is a sequence such that $\rho_{n}=\delta_{2n}$ and
@@
$A_{*,2}^{n}=\{\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq 2n,\,\,\operatorname{Sep}(k_{1},\dots k_{r})=r,\,\,\,\Omega^{k_{j}}_{{\rho}_{n}}\text{ happens}\}.$
We obtain that
@@
$\lim_{n\to\infty}\frac{\mu(A_{*,2}^{n})}{(2n)^{r}\sigma_{\delta_{2n}}^{r}}=1$
using $(M4)_{*}.$ Since $\sigma_{\rho_{n}}=\mu(\Omega_{\rho_{n}})=\mu(\Omega_{\delta_{2n}})=\sigma_{\delta_{2n}},$
$\lim_{n\to\infty}\frac{\mu(A_{*,2}^{n})}{n^{r}\sigma_{\rho_{n}}^{r}}=2^{r}.$
By taking $n=2^{m},$ we have that $\mu(A_{*,2}^{2^{m}})=O\left(2^{rm}{\sigma_{{\rho}_{2^{m}}}^{r}}\right)$ and
$\sum_{m=1}^{\infty}\mu(A_{*,2}^{2^{m}})=O\left({\mathbf{S}}_{r}\right)<\infty.$
@@
It implies that $A_{*,2}^{2^{m}}$ does not occurs for infinitely many $m$’s with probability 1 using the classical Borel-Cantelli Lemma. Then Part (b) follows from the fact that $A_{*}^{n}\subset A_{*,2}^{2^{m}}$ when $2^{m}\leq n\leq 2^{m+1}.$
@@
Part (c) is Theorem 2.12 in [12]. Finally, part (d) holds because in the proof of part (c) in [12], $(M1)_{k}$ and $(M2)_{k}$ for $k=r+1,\dots,2r$ are only used to verify $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$. $\square$
@@
We now turn to this paper’s problem to consider the following targets:
$\Omega_{\rho_{N}}=\{y:\phi(y)>\rho_{N}\}$
and
$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\}$
for $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{t},$ $t\in\mathbb{R}$ and $c_{j}>{\varepsilon}$ for any $j.$ Take $s(N)=K\operatorname{ln}N$ with $K$ being sufficiently large and $\hat{s}(N)=2{\varepsilon}N.$
@@
(a) For any $r\in\mathbb{N},$ $\Omega_{\rho_{N}}$ and $\Omega_{\rho_{N},j}$ satisfy $(M1)_{r}$ and $(M3)_{r}.$
(b) If $x$ is Diophantine then $(M2)_{r}$ holds.
(c) If $x$ is non-periodic then $(M4)_{*}$ and $(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}$ are satisfied.
The proof of Proposition 2.3 is similar to Proposition 3.9 in [12] with some modifications, which is included in Appendix B.
Next, we will recall the exponential mixing property for expanding maps, which is useful not only in the proof Proposition 2.3 but also in the next section. Denote by $BV$ the space of functions with bounded variation on $\mathbb{S}$ and
${\mathbb{E}}\psi=\int\psi{\rm{d}}\mu\,\,\text{for}\,\,\psi\in BV.$
By Proposition 3.8 of [44], there are constants $C,$ $\theta<1$ such that
(2.4) $\left|{\mathbb{E}}\left(\psi_{1}(\psi_{2}\circ f^{n})\right)-({\mathbb{E}}\psi_{1})(E\psi_{2})\right|\leq C\|\psi_{1}\|_{BV}\|\psi_{2}\|_{L^{1}}\theta^{n},\,n\in\mathbb{N}$
for any $\psi_{1},\,\psi_{2}\in BV,$ where $\|\psi\|_{BV}:=\|\psi\|_{L^{1}}+\text{Var}(\psi)$ and $\text{Var}(\psi)$ is the variation of $\psi.$
In the following, we use $C$ to denote a constant that may change from line to line but that will not depend on $N.$
@@
    If $\varphi_{j}\geq 0$ for $1\leq j\leq q,$ we have for any $k_{1}\leq k_{2}\leq\cdots\leq k_{q},$ $k_{i}\in\mathbb{N},$
@@
    (2.6) $\left|{\mathbb{E}}\left(\prod_{j=1}^{q}(\varphi_{j}\circ f^{k_{j}})\right)-\prod_{j=1}^{q}{\mathbb{E}}\varphi_{j}\right|\leq C\theta^{k}\prod_{j=1}^{q}\|\varphi_{j}\|_{BV}.$
@@
where $b_{N}^{\prime}=b_{N}^{\prime\prime}=0$ if $s<1$ and $b_{N}^{\prime},b_{N}^{\prime\prime}$ are chosen to make the expectations of the LHS’s zero if $s>1.$
@@
If $\delta<\alpha$ and either $10<s<$ 1 or $21<s<$ 2 and $x$ is Diophanitne then
[…]
@@
By Lemma 2.5, the limit points of $W_{N}(\cdot)$ and $W_{N}^{\prime}(\cdot)$ are the same. Consider the targets $\Omega_{\rho_{N}}=\{y:\phi(y)>\rho_{N}\}$ for $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\delta}$ and suppose that $x$ is non-periodic. By Proposition 2.3 and Lemma 2.2(b), for large $N$ there are at most $\left[(\delta s)^{-1}\right]$ intervals $I_{j}$ of length $2s(N)$ such that all numbers $n$ such that $X_{n}>N^{1/s}(\operatorname{ln}N)^{\delta}$ are contained in one of those intervals. Moreover taking ${\bar{\alpha}}$ such that $((r+1)s)^{-1}<{\bar{\alpha}}<\alpha,$ we see that among those intervals there are at most $r$ which contain $n^{\prime}$s with $X_{n}\geq N^{1/s}(\operatorname{ln}N)^{{\bar{\alpha}}}.$ We claim that the contribution of the remaining intervals is negligible for a.e. x. Indeed suppose there is an interval $I$ of length $2s(N)$ and ${\bar{n}}\in I$ such that
@@
By the definition of $I$ we have that $Y_{I}\geq{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{\alpha-1}$ for some ${\bar{\varepsilon}}={\bar{\varepsilon}}({\varepsilon})>0.$ Accordingly if the interval $I$ satisfying (2.7), then there are two moments ${\bar{n}},{\hat{n}}\in[1,N]$ such that
@@
Thus the lemma below shows that the contribution of the intervals satisfying (2.7) is negligible. Its proof is available in Appendix B.
@@
For a.e. $y$, (2.8) happens only for finitely many $N.$
@@
To see that all points in ${\mathbb{H}}_{r}$ are realized, fix $r$-tuples $(c_{1},\dots,c_{r})$ and $(t_{1},\dots t_{r})$ such that $c_{j}>0$ for all $j$ and fix $0<{\varepsilon}<c_{j}.$ Consider the events
$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\}.$
Taking $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}$ and applying Proposition 2.3 and Lemma 2.2(d), we see that with probability $1$ there is a subsequence $N_{k}$ and times $n_{1}(k),n_{2}(k),\dots,n_{r}(k)$ such that
$\frac{n_{j}(k)}{N_{k}}\in[t_{j}-{\varepsilon},t_{j}+{\varepsilon}]\quad\text{and}\quad\frac{X_{n_{j}(k)}(y)}{N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}].$
It follows that for each $W\in{\mathbb{H}}_{r}$, $W$ belongs to the limit set with probability $1$ by taking a countable set of ${\varepsilon}_{l}$ converging to $0$ Recall the fact that two elements $W_{1}$ and $W_{2}$ of ${\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$ close if and only if $\{c_{m}^{1}\}_{m=1}^{r}$ is close to $\{c_{m}^{2}\}_{m=1}^{r}$ and $\{t_{m}^{1}\}_{m=1}^{r}$ is close to $\{t_{m}^{2}\}_{m=1}^{r}$. Taking a countable dense subset in ${\mathbb{H}}_{r}$ completes the proof.
@@ 3.1. Splitting the sum
where, as before, $\bar{b}_{N}=\bar{\bar{b}}_{N}=0$ for $s\leq 1$ and they are chosen to make the expectations of LHS’s zero for $s>1.$
@@
[…]
In § 3.2 we consider the case $s<1$ where the argument is much simpler since the sums ${\bar{S}}_{n}$ and ${\bar{{\bar{S}}}}_{n}$ are almost monotone. The case $21\leq s<$ 2 is more involved and requires the use of the reflection principle (see equations (3.14) and (3.23)). In § 3.3 and § 3.4 we prove (3.3) and (3.4) respectively for $21\leq s<$ 2 .
@@
We begin with (3.3). By (2.2) we have
$E\left|\bar{S}_{N}\right|\leq N\int_{0}^{N^{1/s}(\operatorname{ln}N)^{-D}}\frac{a_{s}{\rm{d}}t}{t^{s}}+LN\leq CN^{1/s}(\operatorname{ln}N)^{-D(1-s)}.$
Hence
@@
Take $N_{k}=2^{k}$, then for each $\varepsilon>0$,
@@
where the convergence follows from (3.1). Hence Borel-Cantelli Lemma gives
[…]
Next, consider $2^{k}<N\leq 2^{k+1}.$ Since $|\bar{S}_{N}|\leq|\bar{S}_{N_{k+1}}|+2LN_{k+1}$ and $s>1$, (3.6) gives
[…]
proving (3.3).
Next, we prove (3.4). Choose $D$ such that $D=q\delta$ where $q$ is a positive integer. For $j=0,-1,\cdots,-q,$ let
$X_{k,j}=X_{k}{\mathbf{1}}_{\{X_{k}\in(N^{1/s}(\operatorname{ln}N)^{j\delta},N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}]\}},\quad{\bar{{\bar{S}}}}_{n,j}=\sum_{k=1}^{n}X_{k,j},$
@@
Since $|{\bar{{\bar{S}}}}_{n,j}|\leq|{\bar{{\bar{S}}}}_{N,j}|+2LN$ it suffices to prove that
@@
Next, we claim that for each $m$ there is a constant $C_{m}$ such that
(3.10) $E\left({\bar{{\bar{S}}}}_{N,j}^{m}\right)\leq C_{m}N^{m/s}(\operatorname{ln}N)^{\delta m}.$
Let ${\hat{B}}_{N,j}=\left\{y:\phi(y)\in\left(N^{1/s}(\operatorname{ln}N)^{j\delta},\;N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}\right]\right\},$ $j=0,-1,\cdots,-q.$ Note that
$\frac{{\bar{{\bar{S}}}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\leq\sum_{n=1}^{N}{\mathbf{1}}_{{\hat{B}}_{N,j}}(f^{n}y).$
Thus
(3.11) $E\left(\left(\frac{\bar{\bar{S}}_{N,j}}{N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}}\right)^{m}\right)\leq\sum_{1\leq k_{1}\leq\cdots\leq k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}\left\{f^{k_{j}}y\in{\hat{B}}_{N,j}\right\}\right).$
@@
There is a constant $C$ such that
$\sum_{1\leq k_{1}\leq\cdots\leq k_{m}\leq N}\mu\left(\bigcap_{j=1}^{m}\left\{f^{k_{j}}y\in{\hat{B}}_{N,j}\right\}\right)\leq C\left(N\mu\left({\hat{B}}_{N,j}\right)\right)^{m}.$
Combining the lemma with (2.2) we see that (3.11) is less than $C(\operatorname{ln}N)^{-j\delta sm}.$ Since $j\leq 0$ and $s<1,$ (3.10) follows.
By Markov inequality
$\mu\left(|{\bar{{\bar{S}}}}_{N,j}|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{C}{{\varepsilon}^{m}(\operatorname{ln}N)^{(\alpha-\delta)m}}$
@@
[…]

### v3

@@ 2. An analogue of LIL for Smooth expanding maps
    If $(M4)_{*}$ holds, then with probability 1, the event $A^{N}_{*}$ does not occur for all sufficiently large $N.$
Case 2: ${\mathbf{S}}_{r}=\infty.$
For any $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r}$ and any intervals $I_{1},I_{2}\dots I_{r}\subset[0,1],$
  • (c)
    If $(M1)_{k}$ and $(M2)_{k}$ hold for $k=r,r+1,\dots,2r,$ and $(M3)_{r}$ is satisfied under the condition $\hat{s}(N)/N<\min_{0\leq i<j\leq r}d(I_{i},I_{j})$ where $I_{0}=0,$ then $\mu\left(H_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}\right)=1.$
  • (d)
    If $(M3)_{r}$ is satisfied under the condition $\hat{s}(N)/N<\min_{0\leq i<j\leq r}d(I_{i},I_{j})$ where $I_{0}=0,$ and $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ holds, then $\mu\left(H_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}\right)=1.$
Proof of Lemma 2.2.
@@
Part (a) is Theorem 2.6 in [10]. We now consider the case where ${\mathbf{S}}_{r}<\infty$ and $(M4)_{*}$ holds. Suppose $\{\delta_{N}\}$ is a sequence of positive numbers such that $\rho_{N}=\delta_{2N}.$ Define the event
@@
$A_{*,2}^{N}=\left\{\omega\in\Omega:\,\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq 2N,\,\operatorname{Sep}_{2N}(k_{1},\dots k_{r})=r\,\,\text{such that}\,\,\omega\in\Omega^{k_{j}}_{{\rho}_{N}}\right\}.$
@@
By $(M4)_{*},$ $\limsup_{N\to\infty}\frac{\mu\left(A_{*,2}^{N}\right)}{(2N)^{r}\sigma_{\delta_{2N}}^{r}}\leq C.$ Since $\sigma_{\rho_{N}}\!\!=\!\!\mu\left(\Omega_{\rho_{N}}\right)\!\!=\!\!\mu\left(\Omega_{\delta_{2N}}\right)\!\!=\!\!\sigma_{\delta_{2N}},$ we have
$\limsup_{N\to\infty}\frac{\mu\left(A_{*,2}^{N}\right)}{N^{r}\sigma_{\rho_{N}}^{r}}\leq C.$
By taking $N=2^{M},$ we find that $\mu\left(A_{*,2}^{2^{M}}\right)\leq C2^{rM}{\sigma_{{\rho}_{2^{M}}}^{r}}$ and thus
$\sum_{M=1}^{\infty}\mu\left(A_{*,2}^{2^{M}}\right)\leq C{\mathbf{S}}_{r}<\infty.$
@@
By the classical Borel–Cantelli lemma, $A_{*,2}^{2^{M}}$ does not occurs for infinitely many $M$’s with probability 1. Then Part (b) follows from the fact that $A_{*}^{N}\subset A_{*,2}^{2^{M}}$ when $2^{M}\leq N\leq 2^{M+1}.$
@@
Part (c) is Theorem 2.12 in [10]. Finally, part (d) holds because in the proof of part (c) in [10], $(M1)_{k}$ and $(M2)_{k}$ for $k=r,r+1,\dots,2r$ are only used to verify $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$. ∎
@@
We now address the central problem of this paper by analyzing the following targets:
(2.4) $\Omega_{\rho_{N}}=\left\{y:\phi(y)>C\rho_{N}\right\}$
with $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{t}$ where $t>0$ and $C>0;$
(2.5) $\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\}$
with $c_{j}>0$, $0<{\varepsilon}<c_{j}$ for non-periodic points, and $0<{\varepsilon}<\left(\gamma^{q/s}-1\right)c_{j}/\left(\gamma^{q/s}+1\right)$ for periodic points of period $q$. Also let
(2.6) $\tilde{\Omega}_{\rho_{N},j}=\Omega_{\rho_{N},j}^{c}\mathrel{{$\bigcap$}}f^{-1}\Omega_{\rho_{N},j}^{c}\mathrel{{$\bigcap$}}\cdots\mathrel{{$\bigcap$}}f^{-(p_{0}-1)}\Omega_{\rho_{N},j}^{c}\mathrel{{$\bigcap$}}f^{-p_{0}}\Omega_{\rho_{N},j}$
where $p_{0}=\operatorname{ln}\operatorname{ln}N.$
Additionally, we set $s(N)=K\operatorname{ln}N$, where $K$ is chosen to be sufficiently large, and set $\hat{s}(N)=2\varepsilon N$.
@@
  • (a)
    The events $\left\{\Omega_{\rho_{N}}\right\}$ and $\left\{\Omega_{\rho_{N},j}\right\}$ satisfy conditions $(M1)_{r}$ and $(M3)_{r}$ for all $r\in\mathbb{N}.$
  • (b)
    If $x$ is Diophantine then condition $(M2)_{r}$ holds for any $r\in\mathbb{N}.$
  • (c)
    Condition $(M4)_{*}$ is satisfied for $\left\{\Omega_{\rho_{N}}\right\}.$
  • (d)
    Suppose that $\min_{i\neq j}d(I_{i},I_{j})>{\varepsilon}$ for some ${\varepsilon}>0.$ If $x$ is not a periodic point, condition $(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}$ holds for the events $\left\{\Omega_{\rho_{N},j}\right\}.$ Moreover, if $x$ is periodic, the same condition holds when $\left\{\Omega_{\rho_{N},j}\right\}$ is replaced by the events $\left\{\tilde{\Omega}_{\rho_{N},j}\right\}$ in the definition of $A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}.$
The proof of Proposition 2.3 follows a similar structure to that of Proposition 3.9 in [10] with some modifications, which is provided in Appendix B.
Next, we will recall the exponential mixing property for expanding maps, which will be useful not only in the proof Proposition 2.3 but also in the next section. We denote $BV$ to be the space of functions with bounded variation on $\mathbb{S},$ and ${\mathbb{E}}\psi=\int\psi{\rm{d}}\mu\,\,\,\text{for}\,\,\,\psi\in L^{1}(\mu).$ By Proposition 3.8 of [31], there exist constants $C>0,$ $\theta<1$ such that for all $\psi_{1},\,\psi_{2}\in BV,$
(2.7) $\left|{\mathbb{E}}\left(\psi_{1}(\psi_{2}\circ f^{n})\right)-({\mathbb{E}}\psi_{1})(E\psi_{2})\right|\leq C\|\psi_{1}\|_{BV}\|\psi_{2}\|_{L^{1}}\theta^{n},\,n\in\mathbb{N},$
where $\|\psi\|_{BV}:=\|\psi\|_{L^{1}}+\text{Var}(\psi),$ and $\text{Var}(\psi)$ denotes the variation of $\psi.$
@@
    If $\varphi_{j}\geq 0$ for $1\leq j\leq q,$ then for any $k_{1}\leq k_{2}\leq\cdots\leq k_{q}$ with $k_{i}\in\mathbb{N},$ we have
@@
    (2.9) $\left|{\mathbb{E}}\left(\prod_{j=1}^{q}\varphi_{j}\circ f^{k_{j}}\right)-\prod_{j=1}^{q}{\mathbb{E}}\varphi_{j}\right|\leq C\theta^{k}\prod_{j=1}^{q}\|\varphi_{j}\|_{BV}.$
@@
where $b_{N}^{\prime}=b_{N}^{\prime\prime}=0$ if $s\leq 1$ and $b_{N}^{\prime},b_{N}^{\prime\prime}$ are chosen to so that the expectations of the left-hand sides are zero when $s>1.$
@@
If $\delta<\alpha$ and either $0<s<1$ or $1\leq s<2$ and $x$ is Diophanitne then
$\max_{1\leq n\leq N}\frac{\left|S_{n}^{\prime\prime}\right|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\rightarrow 0\quad\text{a.s.}.$
@@
By Lemma 2.5, the limit points of $W_{N}(\cdot)$ and $W_{N}^{\prime}(\cdot)$ are the same. Consider the targets
$\Omega_{\rho_{N}}=\{y:\phi(y)>\rho_{N}\},$
where $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\delta}.$ By Proposition 2.3(c) and Lemma 2.2(b), for sufficiently for large $N,$ there exist at most $\left[(\delta s)^{-1}\right]$ intervals $I_{j}$ of length $2s(N)$ such that every $n$ satisfying $X_{n}>N^{1/s}(\operatorname{ln}N)^{\delta}$ is contained in one of those intervals.
Moreover, by choosing $\bar{\alpha}$} such that $((r+1)s)^{-1}<{\bar{\alpha}}<\alpha,$ we see that among those intervals, at most $r$ contain $n$ for which
$X_{n}\geq N^{1/s}(\operatorname{ln}N)^{{\bar{\alpha}}}.$
We claim that the contribution of the remaining intervals is negligible for almost every $y.$ Indeed, suppose there is an interval $I$ of length $2s(N)$ and ${\bar{n}}\in I$ such that
@@
By the definition of $I,$ we have
$Y_{I}\geq{\bar{\varepsilon}}N^{1/s}(\operatorname{ln}N)^{\alpha-1}$
for some ${\bar{\varepsilon}}={\bar{\varepsilon}}({\varepsilon})>0.$ Accordingly, if the interval $I$ satisfies (2.10), then there exist two moments ${\bar{n}},{\hat{n}}\in[1,N]$ such that
@@
Thus, the following lemma establishes that the contribution of the intervals satisfying (2.10) is negligible. Its proof is provided in Appendix B.
@@
For almost every $y,$ (2.11) happens only for finitely many $N.$
@@
To verify that all points in ${\mathbb{H}}_{r}$ are realizable, we fix an $r$-tuple $(c_{1},\dots,c_{r})$ such that $c_{j}>0$ for all $j,$ together with an increasing $r$-tuple $(t_{1},\dots,t_{r})$ satisfying $t_{1}<t_{2}<\cdots<t_{r}.$
If $x$ is not a periodic point, we define the events
$\Omega_{\rho_{N},j}=\left\{y:\frac{\phi(y)}{\rho_{N}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}]\right\},\,\,1\leq j\leq r$
for $0<{\varepsilon}<c_{j}.$
If $x$ is a periodic point of period $q,$ we instead choose $0<{\varepsilon}<\left(\gamma^{q/s}-1\right)c_{j}/\left(\gamma^{q/s}+1\right)$ and consider the events
$\tilde{\Omega}_{\rho_{N},j}=\Omega_{\rho_{N},j}^{c}\mathrel{{\bigcap}}f^{-1}\Omega_{\rho_{N},j}^{c}\mathrel{{\bigcap}}\cdots\mathrel{{\bigcap}}f^{-(p_{0}-1)}\Omega_{\rho_{N},j}^{c}\mathrel{{\bigcap}}f^{-p_{0}}\Omega_{\rho_{N},j},\,\,1\leq j\leq r,$
where $p_{0}=\operatorname{ln}\operatorname{ln}N.$
Taking $\rho_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}$ and applying Proposition 2.3 (d) and Lemma 2.2 (d), we conclude that with probability $1$ there is a subsequence $N_{k}$ and times $n_{1}(k),n_{2}(k),\dots,n_{r}(k)$ such that
$\frac{n_{j}(k)}{N_{k}}\in[t_{j}-2{\varepsilon},t_{j}+2{\varepsilon}]\quad\text{and}\quad\frac{X_{n_{j}(k)}(y)}{N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}}\in[c_{j}-{\varepsilon},c_{j}+{\varepsilon}].$
It follows that for each $\in{\mathbb{H}}_{r}, WW$ belongs to the limit set with probability $1$ by taking a countable set of ${\varepsilon}_{l}$ converging to $0.$ Recall the fact that two elements $W_{1}$ and $W_{2}$ of ${\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$ close if and only if $\{c_{m}^{1}\}_{m=1}^{r}$ is close to $\{c_{m}^{2}\}_{m=1}^{r}$ and $\{t_{m}^{1}\}_{m=1}^{r}$ is close to $\{t_{m}^{2}\}_{m=1}^{r}$. Taking a countable dense subset in ${\mathbb{H}}_{r}$ completes the proof.
@@ 3.1. Splitting the sum
where, as before, $\bar{b}_{N}=\bar{\bar{b}}_{N}=0$ for $s\leq 1$ and they are chosen so that the expectations of the left-hand sides are zero for $s>1.$
@@
(3.4) $\max_{1\leq n\leq N}\frac{|{\bar{{\bar{S}}}}_{n}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\to 0\;\;\text{a.s.}.$
In § 3.2, we consider the case $s<1,$ where the argument is much simpler, since the sums ${\bar{S}}_{n}$ and ${\bar{{\bar{S}}}}_{n}$ are almost monotone. The case $1\leq s<2$ is more involved and requires the application of the reflection principle (see equations (3.14) and (3.23)). In § 3.3 and § 3.4, we establish (3.3) and (3.4), respectively, for $1\leq s<2.$
@@
We begin by considering (3.3). By (2.1), we have
$E\left|\bar{S}_{N}\right|\leq N\int_{0}^{N^{1/s}(\operatorname{ln}N)^{-D}}\frac{a_{s}{\rm{d}}t}{t^{s}}+LN\leq CN^{1/s}(\operatorname{ln}N)^{-D(1-s)},$
for $L=\max_{x\in\mathbb{S}}\left|\psi(x)\right|.$ Hence
@@
Let $N_{k}=2^{k}.$ For each $\varepsilon>0$, we have
@@
where the convergence follows from (3.1). By the classical Borel-Cantelli lemma
(3.6) $\frac{|\bar{S}_{N_{k}}|}{N_{k}^{1/s}(\operatorname{ln}N_{k})^{\alpha}}\rightarrow 0\quad\text{a.s.}.$
Next, consider the case $2^{k}<N\leq 2^{k+1}.$ Since $|\bar{S}_{N}|\leq|\bar{S}_{N_{k+1}}|+2LN_{k+1}$ and $s<1$, (3.6) yields
(3.7) $\frac{|\bar{S}_{N}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\rightarrow 0\quad\text{ a.s.}$
which completes the proof of (3.3).
$\mu({\hat{B}}_{N,j})\sim 2a^{s}h(x)N^{-1}(\operatorname{ln}N)^{-j\delta s}.$
Next, we prove (3.4). Let $D$ be chosen such that $D=q\delta,$ where $q$ is a positive integer. For $j=0,-1,\cdots,-q,$ define
(3.8) ${\hat{B}}_{N,j}=\left\{y:\phi(y)\in\left(N^{1/s}(\operatorname{ln}N)^{j\delta},\;N^{1/s}(\operatorname{ln}N)^{(j+1)\delta}\right]\right\},$
and
(3.9) ${\bar{{\bar{S}}}}_{n,j}=\sum_{k=1}^{n}\phi(f^{k}y){\mathbf{1}}_{{\hat{B}}_{N,j}}(f^{k}y),$
@@
Since $|{\bar{{\bar{S}}}}_{n,j}|\leq|{\bar{{\bar{S}}}}_{N,j}|+2LN$ and $s<1,$ it suffices to prove that
@@
  • (a)
    If $x$ is not a periodic point then ${\mathbb{E}}{\bar{{\bar{S}}}}_{N,j}^{m}\leq CN^{m/s}(\operatorname{ln}N)^{\delta m}.$
  • (b)
    If $x$ is periodic then ${\mathbb{E}}{\bar{{\bar{S}}}}_{N,j}^{m}\leq CN^{m/s}(\operatorname{ln}N)^{\delta m}\left(\operatorname{ln}\operatorname{ln}N\right)^{m}.$
By Markov’s inequality
$\mu\left(|{\bar{{\bar{S}}}}_{N,j}|\geq{\varepsilon}N^{1/s}(\operatorname{ln}N)^{\alpha}\right)\leq\frac{C\left(\operatorname{ln}\operatorname{ln}N\right)^{m}}{{\varepsilon}^{m}(\operatorname{ln}N)^{(\alpha-\delta)m}}$
@@
(3.12) $\frac{|\bar{\bar{S}}_{N,j}|}{N^{1/s}(\operatorname{ln}N)^{\alpha}}\rightarrow 0\quad\text{a.s.}.$